A student dissolves 5.0 g NaOH in 100 mL of water in a calorimeter at 20.5°C. As the NaOH dissolves, the temperature s of the surrounding water rises to 31.00°C. Determine the heat of solution of NaOH in J/g.


1
Expert's answer
2021-03-05T04:06:53-0500

Mass of NaOH = 5g

Mass of Water = density x volume = 1x100 = 100g

Temperature change= (31.00 - 20.5) = 10.5°C

Specific heat capacity of water = 4.18


Heat absorbed by the calorimeter = (5+100) x 4.18 x 10.5 =4608.45J


Since the reaction is exothermic, thermal energy evolved in the process = -q cal = -4608.45J



Heat of solution of NaOH = -4608.45J/5.0g = -921.69J/g


Heat of solution of NaOH is = -921.69J/g